# Learning Quantum Processes with Quantum Statistical Queries

### v1

@@ 6 Application of QPSQ in Cryptanalysis of CR-QPUFs
In this section, we present applications of our QPSQ model in the realm of cryptanalysis. Specifically, we focus on a particular class of PUFs known as Classical Readout Quantum Physically Unclonable Functions (CR-QPUFs). Our objective is to investigate how our proposed algorithm can serve as an effective strategy for attacking their security. We have chosen this specific cryptographic primitive as our case study due to the inherent compatibility between the conceptual and formal framework of CR-QPUFs and the QPSQ framework. Our results showcase a concrete application of a particular learning model and its associated algorithms in the field of cryptography, thereby establishing a connection between the domains of quantum learning theory and quantum cryptography. In the next subsection, we will delve into our new attacks and the results and we will see that these outcomes not only address an ongoing open question regarding the security of different classes of quantum physically unclonable functions but also bridge the existing gap in the security analysis of CR-QPUFs, by demonstrating the unlikelihood of achieving the desired level of security for this specific class against efficient quantum adversaries.
@@
In this section, we attempt to provide an almost definitive and theoretical answer to this question, made possible through the application of our QPSQ framework and the algorithms we have devised. We apply these tools to various iterations of CR-QPUFs. Through this in-depth case study, we conclusively demonstrate that no matter the complexity or depth of the quantum circuit associated with CR-QPUFs, they remain susceptible to learning (within the relevant learning model we will elaborate upon), hence we cannot hope to achieve a secure protocol relying on them. The selection of CR-QPUFs as an ideal subject for our model comes from multiple reasons. Notably, these PUFs have been formally defined within the statistical query model, as noted in [69]. Our investigation reveals that the susceptibility of this type of PUF fundamentally arises from the fact that a large class of quantum processes can be efficiently learned in the QPSQ model. This spectrum includes processes such as Haar random unitaries which are random and extremely complicated to construct when scaling with the system size. Consequently, one cannot hope to evade the learning attacks. This finding, counterintuitive and surprising, highlights a critical fact regarding the security of the cryptographic schemes surrounding hardware assumptions, like physical unclonability: Security does not inherently rely on process complexity alone or the presumption that the quantum process cannot be fully and efficiently learnt by tomography. Instead, the crucial factors are the verification method and the nature of the data accessible to the adversary, which bears an important message for the design of secure protocols and subroutines based on hardware and physical assumptions.
Before we dive into the result, we give a brief overview of the desired security property associated with the physically unclonable function and their main application, namely authentication, and then we present a structure for the authentication protocols based on CR-QPUFs, which allows us to better demonstrate our cryptanalysis using QPSQs and show the effectiveness of our attacks.
@@ 6.1 Security property and attack model
As discussed in [6], the main cryptographic notion associated with PUFs in general, and particularly for the use-case of authentication and identification, is unforgeability. Essentially, unforgeability means that an adversary having access to an efficient-size database of input-output pairs (also called challenge-response pairs or CRPs in the context of PUF) of a specific function of interest, should not be able to reproduce the output of the function on a new input. This database in many cases is obtained via the adversary querying an oracle that realises the black-box access to the function. The unforgeability in cryptography is often defined as a formal game between a challenger and an adversary, and the forgery attacks are considered as algorithms (performed by the adversary) that win this game with non-negligible probability. Unforgeability is also extended to the quantum world generally in two directions: It can also be defined for quantum processes (instead of classical functions) [29] and for quantum adversaries having access to quantum oracles of a classical function [16, 3, 29]. As there are many levels and flavours of unforgeability that involve a quantum adversary, and the technical discussions about different formal definitions of it are outside the scope of this paper, we avoid analysing the security through formal unforgeability games and instead, we demonstrate our attacks using the security of the protocol. However, we note that the security inherently relies on this cryptographic notion which has a close connection to the notion of learnability, especially when considering quantum processes.
@@ 6.2 The general structure of a CR-QPUF-based authentication protocol
Given the adversarial model in the previous section, we now define a general authentication protocol for CR-QPUFs and discuss how their security (or soundness) is defined. Later we go over specific instances of such protocols for different choices of CR-QPUFs, distributions of the challenge states and observables. For the CR-QPUFs, we work with a definition similar to [69] based on statistical queries. However, in order to use our learning attacks, we need to adapt the definition to our framework. We will discuss how our framework easily and naturally generalises to this case. However, for now, we denote the CR-QPUFs as $\mathcal{C}$ and abstractly define it as a process, associated with the completely positive trace preserving (CPTP) map $\mathcal{E}$ over the $n-$qubit state space, able to produce statistical queries for any challenge in the challenge set, given an observable $O$ and a tolerance parameter $\tau$.
@@
      if t=c: $\mathcal{P}$ creates $poly(1/\tau)$ copies of the state $\rho_{i}=\rho(x_{i})$.
@@
We say the CR-QPUF-based authentication protocol defined as 3 is secure if the success probability of any QPT adaptive adversary $\mathcal{A}$ in producing an output $\tilde{y}$ for any $x$ sampled at random from a database $D$, over a distribution $\mathcal{D}$, that passes the verification i.e. satisfies the condition $|y-\tilde{y}|\leq 2\tau$, is at most negligible in the security parameter.
@@ 6.3 CR-QPUFs in QPSQ framework
While we have been discussing the QPStat oracles abstractly so far, the CR-QPUF device must, in practice, be able to take a quantum state, an observable and a tolerance parameter $\tau$, and output with high-probability a $\tau$-estimate of the expectation value. When a query is made to ${}_{\mathcal{E}}$, the device would apply $\mathcal{E}$ to multiple copies of the state, estimate the expectation value of the observable and respond with that value as output. There are multiple methods for this estimation, such as those shown in [48, 41, 49, 85], usually requiring measurements on $poly(1/\tau)$ copies of the state for statistical estimation. In a real implementation of the protocol, the device would receive copies of the input state, either through a quantum communication channel or prepared by the device given a classical description as input.
@@
For a challenge unitary $U_{in}$, the response $\hat{R}_{out}$ is the mean of a set of i.i.d. samples from running $U_{in}$ on the noisy quantum device identified by $qPUF_{id}$ and measuring in the computational basis. Corresponding to the uniquely noisy device $id$ and unitary challenge $U_{in}$, we have the quantum process $\Lambda_{in}^{id}$. This gives rise to the quantum state $\rho_{in}^{id}=\Lambda_{in}^{id}(|0\rangle\langle 0|)$. Let the associated Born distribution be $\mathcal{P}_{U_{in},id}(x)=tr(M_{x}\rho_{in}^{id}M_{x}^{\dagger})$, where $M_{x}$ is the measurement operator corresponding to outcome $x$. The output of the statistical query oracle $QEval$ associated with some efficiently computable function $\phi\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ is then given by
$\hat{R}_{out}\leftarrow QEval(qPUF_{id},U_{in}),\text{ }|\mathbb{E}_{x\sim\mathcal{P}_{U_{in},id}}[\phi(x)]-\hat{R}_{out}|\leq\tau$ (32)
@@
6.4 Cryptanalysis results
@@
Algorithm 4 Attack on Protocol 3 in QSQ with observable $O$, tolerance $\tau$
@@
Consider the attack described in Algorithm 4. From equation 14, we can see that the adversary makes $poly(1/\tau)$ statistical queries of tolerance $\tau$, and the algorithm runs in $poly(1/\tau)$ time. Algorithm 4 is thus an efficient QPSQ learning algorithm.
@@
7 Conclusion and Future Works
@@
One of the main advantages of our QPSQ model is that due to its generality, it allows for the formal study of various algorithms for quantum process learning which may have been designed within different disciplines and approaches. Our arbitrary process learning algorithm imposes no restrictions on the process itself but guarantees performance only for local observables and over locally flat distributions. An exciting avenue for future research is to expand this study further and seek or design algorithms with fewer restrictions on quantum process learning.
For Algorithm 2, our observations indicate that, in some applications, overcoming the locally flat distribution restriction would open up a wide range of possibilities, even if it meant restricting the family of predictable quantum processes. Conversely, in other applications, the ability to learn the process with non-local observables would be of paramount importance. The redesign of our algorithm with these alternative sets of assumptions represents an exciting area for future work, with implications for the applications of the equivalent shadow-based algorithm presented in [46].
In terms of applications, we believe there is much to explore in both proposed directions: cryptanalysis and learning theory. First, in the case of the cryptanalysis of CR-QPUFs, our attack formally breaks the security of a large class of them, for a wide range of observables. However, as discussed in section 6.2, there exists a class of efficiently estimable observables that are non-local, falling outside the formal guarantee of our attack. Therefore, while our work addresses a significant portion of the security analysis for these primitives, a narrow gap remains. It remains an open question whether it’s possible to design an efficient authentication protocol based on CR-QPUFs, using such non-local observables while maintaining resistance to learning attacks. Nonetheless, it’s worth noting that, given their definition within the statistical query model and the results presented in this paper, even if such secure protocols exist, proving their security against the powerful learning algorithms for QPSQ would be a challenging task.
@@
Another interesting future direction is the examination of cryptographic primitives and schemes beyond CR-QPUFs through the lens of learning quantum processes with quantum statistical queries. Investigating the cryptographic implications of the efficient learners in our model as adversarial strategies presents an intriguing direction for further study.
In the realm of learning theory, our newly proposed QPSQ model provides a natural framework to study the learnability of quantum processes using statistical queries. Potential research directions using this framework include not only the study of classical SQ learning results in a quantum setting but also the adaptation of classical techniques to this setting. One such direction would be to adapt classical techniques on establishing lower bounds for learning concept classes, to learning processes in the QPSQ setting. Another compelling question in this context is whether we can show a formal separation between QPSQ and classical learners, indicating a potential path towards finding quantum advantage in learning problems.
@@
Furthermore, it would be intriguing to observe the implementation of our learner on actual hardware or its application to data acquired from real physical experiments.
@@
Acknowledgements: The authors thank Armando Angrisani for his valuable inputs and discussions, especially the discussions towards establishing the definition of QPSQ, and for sharing related results with us during the project. We also thank Elham Kashefi, Dominik Leichtle, Yao Ma, and Sean Thrasher for interesting discussions and comments at different stages of this project. The authors acknowledge the support of the Quantum Advantage Pathfinder (QAP), with grant reference EP/X026167/1 and the UK Engineering and Physical Sciences Research Council.

### v3

@@ 7 Characterization of CR-QPUF Security
In this section, we present applications of our QPSQ model in the realm of cryptography. Specifically, we focus on a particular class of PUFs known as Classical Readout Quantum Physically Unclonable Functions (CR-QPUFs). Our objective is to investigate how our proposed algorithm can serve as an effective strategy for attacking their security, as well as to use our hardness results to study the possibility of a secure construction. We have chosen this specific cryptographic primitive as our case study due to its inherent compatibility with the conceptual and formal QPSQ framework. Our results showcase a concrete application of a particular learning model and its associated algorithms in the field of cryptography, thereby establishing a connection between the domains of quantum learning theory and quantum cryptography. In the next subsection, we will delve into our new results and we will see that these outcomes not only address an ongoing open question regarding the security of different classes of quantum physically unclonable functions but also significantly narrow the existing gap in the security analysis of CR-QPUFs, by demonstrating broad conditions under which this class of protocols is vulnerable against efficient quantum adversaries.
@@
In this section, we attempt to provide an almost definitive and theoretical answer to this question, made possible through the application of our QPSQ framework, the learning algorithm we have devised and our hardness results for learning certain classes of unitaries. We apply these tools to various iterations of CR-QPUFs. Through this in-depth case study, we show that for most known practical scenarios CR-QPUFs remain susceptible to learning attacks irrespective of the complexity or depth of the associated quantum circuit, and we demonstrate precisely the conditions under which these learning attacks (within the relevant learning model we will elaborate upon), are guaranteed to succeed efficiently. The selection of CR-QPUFs as an ideal subject for our model comes from multiple reasons. Notably, these PUFs have been formally defined within the statistical query model, as noted in [52]. Our investigation reveals that the susceptibility of this type of PUF fundamentally arises from the fact that a large class of quantum processes can be efficiently learned in the QPSQ model when the measurement performed is local. This spectrum includes processes such as Haar random unitaries which are random and extremely complicated to construct when scaling with the system size. Consequently, one cannot hope to evade the learning attacks in the local measurement regime. This finding highlights a critical fact regarding the security of the cryptographic schemes surrounding hardware assumptions, like physical unclonability: Security does not inherently rely on process complexity alone or the presumption that the quantum process cannot be fully and efficiently learnt by tomography. Instead, the crucial factors are the verification method and the nature of the data accessible to the adversary, which bears an important message for the design of secure protocols and subroutines based on hardware and physical assumptions, as well as more generally, cases where the adversary is assumed certain quantum oracle access to the cryptographic primitive.
Before we dive into the result, we give a brief overview of the desired security property associated with the physically unclonable function and their main application, namely authentication, and then we present a structure for the authentication protocols based on CR-QPUFs, showing how they fit into the QPSQ learning model. This allows us to analyze this class of protocols using our new learning results directly.
@@ 7.1 Security property and attack model
As discussed in [50], the main cryptographic notion associated with PUFs in general, and particularly for the use-case of authentication and identification, is unforgeability. Essentially, unforgeability means that an adversary having access to an efficient-size database of input-output pairs (also called challenge-response pairs or CRPs in the context of PUF) of a specific function of interest, should not be able to reproduce the output of the function on a new input. This database in many cases is obtained via the adversary querying an oracle that realises the black-box access to the function. The unforgeability in cryptography is often defined as a formal game between a challenger and an adversary, and the forgery attacks are considered as algorithms (performed by the adversary) that win this game with non-negligible probability. Unforgeability is also extended to the quantum world generally in two directions: It can also be defined for quantum processes (instead of classical functions) [94] and for quantum adversaries having access to quantum oracles of a classical function [46, 95, 94]. As there are many levels and flavours of unforgeability that involve a quantum adversary, and the technical discussions about different formal definitions of it are outside the scope of this paper, we avoid analysing the security through formal unforgeability games and instead, we study the security of the protocol through the lens of learnability. However, we note that the security inherently relies on this cryptographic notion which has a close connection to the notion of learnability, especially when considering quantum processes.
@@ 7.2 The general structure of a CR-QPUF-based authentication protocol
Given the adversarial model in the previous section, we now define a general authentication protocol for CR-QPUFs and discuss how their security (or soundness) is defined. Later we go over specific instances of such protocols for different choices of CR-QPUFs, distributions of the challenge states and observables. For the CR-QPUFs, we work with a definition similar to [52] based on statistical queries. However, in order to use our learning results, we need to adapt the definition to our framework. We will discuss how our framework easily and naturally generalises to this case. However, for now, we denote the CR-QPUFs as $\mathcal{C}$ and abstractly define it as a process, associated with the completely positive trace preserving (CPTP) map $\mathcal{E}$ over the $n$-qubit state space, able to produce statistical queries for any challenge in the challenge set, given an observable $O$ and a tolerance parameter $\tau$.
@@
      if t=c: $\mathcal{P}$ creates $\mathsf{poly}(1/\tau)$ copies of the state $\rho_{i}=\rho(x_{i})$.
@@
We say the CR-QPUF-based authentication protocol 3 is secure if the success probability of any QPT adaptive adversary $\mathcal{A}$ in producing an output $\tilde{y}$ for any $x$ sampled at random from a database $D$, over a distribution $\mathcal{D}$, that passes the verification i.e. satisfies the condition $|y-\tilde{y}|\leq 2\tau$, is at most negligible in the security parameter.
@@ 7.3 CR-QPUFs in QPSQ framework
While we have been discussing the QPStat oracles abstractly so far, the CR-QPUF device must, in practice, be able to take a quantum state, an observable and a tolerance parameter $\tau$, and output with high-probability a $\tau$-estimate of the expectation value. When a query is made to QPStatE, the device would apply $\mathcal{E}$ to multiple copies of the state, estimate the expectation value of the observable and respond with that value as output. There are multiple methods for this estimation, such as those shown in [32, 74, 75, 76], usually requiring measurements on $\mathsf{poly}(1/\tau)$ copies of the state for statistical estimation. In a real implementation of the protocol, the device would receive copies of the input state, either through a quantum communication channel or prepared by the device given a classical description as input.
@@
For a challenge unitary $U_{in}$, the response $\hat{R}_{out}$ is the mean of a set of i.i.d. samples from running $U_{in}$ on the noisy quantum device identified by $qPUF_{id}$ and measuring in the computational basis. Corresponding to the uniquely noisy device $id$ and unitary challenge $U_{in}$, we have the quantum process $\Lambda_{in}^{id}$. This gives rise to the quantum state $\rho_{in}^{id}=\Lambda_{in}^{id}(|0\rangle\langle 0|)$. Let the associated Born distribution be $\mathcal{P}_{U_{in},id}(x)=Tr(M_{x}\rho_{in}^{id}M_{x}^{\dagger})$, where $M_{x}$ is the measurement operator corresponding to outcome $x$. The output of the statistical query oracle $QEval$ associated with some efficiently computable function $\phi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ is then given by
$\hat{R}_{out}\leftarrow QEval(qPUF_{id},U_{in}),\text{ }|\mathbf{E}_{x\sim\mathcal{P}_{U_{in},id}}[\phi(x)]-\hat{R}_{out}|\leq\tau$ (108)
@@
7.4 Vulnerability from learning results
@@
Algorithm 4 QPSQ attack on Protocol 3 with observable $O$, tolerance $\tau$
@@
Consider the attack described in Algorithm 4. From equation 18, we can see that the adversary makes $\mathsf{poly}(1/\tau)$ statistical queries of tolerance $\tau$, and the algorithm runs in $\mathsf{poly}(1/\tau)$ time. Algorithm 4 is thus an efficient QPSQ learning algorithm.
@@
Remark 4.
Naively, one might anticipate that our hardness of learning results would imply a positive security result for CR-QPUFs. However, due to the nature of the verification procedure of the protocol, an adversary only needs to learn to predict approximately correct expectation values, rather than learn the underlying process up to accuracy in diamond distance. This is a significantly easier task, and as such, the hardness results cannot directly be used to prove the security of this protocol. In fact, contrary to this anticipation, we observe the opposite outcome. When the underlying process of the CR-QPUF is chosen from a hard-to-learn ensemble, such as a (approximate) unitary 2-design, then the output expectation values are highly concentrated regardless of the choice of observable and input state, enabling an easy attack with high success probability for multiple rounds. This fact, along with Theorem 6 shows that most efficiently implementable CR-QPUFs do not satisfy soundness against adaptive adversaries.
8 Conclusion and Future Work
@@
One of the main advantages of our QPSQ model is that due to its generality, it allows for the formal study of various algorithms for quantum process learning which may have been designed within different disciplines and approaches. Algorithm 2 for learning arbitrary processes imposes no restrictions on the process itself but guarantees performance only for local observables and over locally flat distributions. An exciting avenue for future research is to expand this study further and seek or design algorithms with fewer restrictions on quantum process learning. Our observations indicate that, in some applications, overcoming the locally flat distribution restriction would open up a wide range of possibilities, even if it meant restricting the family of predictable quantum processes. Conversely, in other applications, the ability to learn the process with non-local observables would be of paramount importance. Redesigning the algorithm with these alternative sets of assumptions represents an exciting area for future work, with implications for the applications of the equivalent shadow-based algorithm presented in [53].
We have shown the hardness of learning exact and approximate 2-designs and Haar-random unitaries in our statistical query model. In the case of Haar-random unitaries, the bound obtained in [60] for learning the output distributions of Haar-random circuits (from a weaker access model) is doubly exponential, making it stronger than ours by an exponential factor. It is an interesting open question whether we can obtain a matching bound in our stronger access model. Following in the direction of [60], it would also be interesting to see what lower bound can be obtained for constant and logarithmic depth circuits in our model.
Furthermore, it is interesting to compare learnability in this framework to the usual quantum statistical query framework for learning states. When considering the task of learning classical functions encoded either as a unitary or as a quantum example, it is interesting to see whether the additional choice of input state available to a learner in our model can provide any advantage, i.e. are there any separations between QPSQ and QSQ learners when looking at classical functions?. Another compelling question in this context is whether we can show a meaningful, formal separation between QPSQ and classical learners (beyond what has already been shown through generalizing QSQs) indicating a potential path towards finding quantum advantage in learning problems.
In terms of applications, there is much to explore in cryptography. While we have shown that efficient CR-QPUFs cannot be secure, we have not been able to conclusively prove that no CR-QPUF protocol can satisfy soundness. This poses an interesting open question for future work.
@@
Another interesting future direction is the examination of cryptographic primitives and schemes beyond CR-QPUFs through the lens of learning quantum processes with quantum statistical queries. Investigating the cryptographic implications of learning algorithms in our model as adversarial strategies presents an intriguing direction for further study. Another interesting direction is to identify cryptographic schemes whose security can be proven from the application of our lower bounds in the QPSQ model.
@@
Finally, it would be intriguing to observe the implementation of our learner on actual hardware or its application to data acquired from real physical experiments.
@@
Acknowledgements: The authors thank Armando Angrisani for his valuable inputs and discussions, especially the discussions towards establishing the definition of QPSQ, and for sharing related results from [59] with us during the project. We thank Alexander Nietner for pointing us towards the extension of Lemma 11 to Theorem 5, as well as noticing an error in a security result regarding CR-QPUFs in a previous version of this work. We thank Eliott Mamon for regularly participating in our discussions and verifying some of our mathematical proofs. We also thank Elham Kashefi, Dominik Leichtle, Yao Ma, and Sean Thrasher for interesting discussions and comments at different stages of this project. The authors acknowledge the support of the Quantum Advantage Pathfinder (QAP), with grant reference EP/X026167/1 and the UK Engineering and Physical Sciences Research Council.